The spot EUR/USD is 1.12 and the forward rate is 1.1. The interest rate in France is 3% and 4% in the US. a) Does the iRP hold? b) If not, how could you make a CIA profit by using 1000 EUR? Show your work. c) What is the forward rate that would make CIA disappear?
Answer #1

Acc to IRPT

Fwd Rate = Spot Rate * ( 1+Hi) / (1+Fi)

Hi - Int Rate in US

Fi = Int Rate in Europe

Fwd Rate = Spot Rate * ( 1+Hi) / (1+Fi)

= $ 1.12 * ( 1+ 0.04) / ( 1 +0.03 )

= $ 1.12 * ( 1.04) / ( 1.03 )

= $ 1.1309

Actual Fwd rate is $ 1.1 Thus IRPT doesn' holds good. And there is chance for CIA.

Part 2:

CIA with Euro 1000:

Take Loan of Euro 1000

COnvert into USD using spot Rate = 1000 * 1.12

= 1120

Invest in USD for 1 year and Realize after 1 Year:

= 1120 * 1.04

= 1164.8

Convert the realized amount into Euro using FWD Rate = 1164.8 / 1.1

= Eur 1058.91

Clear the loan with Int = 1000 * 1.03

= 1030

Book Profit = 1058.91 - 1030

= Eur 28.91

Part C:

Acc to IRPT

Fwd Rate = Spot Rate * ( 1+Hi) / (1+Fi)

Hi - Int Rate in US

Fi = Int Rate in Europe

Fwd Rate = Spot Rate * ( 1+Hi) / (1+Fi)

= $ 1.12 * ( 1+ 0.04) / ( 1 +0.03 )

= $ 1.12 * ( 1.04) / ( 1.03 )

= $ 1.1309

if Fwd rate is 1 EUR = $ 1.1309, IRPT holds good and CIA will disappear.